# Rodrigues Network for Learning Robot Actions

### v1

@@ Abstract
**footnotetext: Equal contribution\$$\ddagger$\$$\ddagger$\$footnotetext: Work performed while visiting Stanford University.
@@ 1 Introduction
We study the problem of understanding and predicting the actions of articulated actors. Articulated actors (e.g. robots [1] and animated characters [40, 31]) are systems that use multiple rotational joints to connect moving links. Their actions, including poses, motions, or control commands, are usually represented as values associated with all joints. Learning with articulated actors usually involves predicting their actions while processing diverse sensory inputs. Such a problem lies in a wide spectrum of intelligent systems, from whole-body controllers [34, 27], to dexterous grasp detectors [18, 53, 50, 56], to motion retargetting networks [2, 20].
@@
To learn the actions of articulated actors, we seek a neural network architecture that can both understand robot structures and enforce structural constraints in its predictions. However, most existing network architectures, such as MLPs and Transformers [48], treat actions as unstructured tokens, ignoring the kinematic relationships that govern joint interactions. A few prior works have incorporated structural priors like robot link connectivity, using graph convolutions [54, 16, 2] or masked attention [41]. Yet, the design of network architectures that fully capture and enforce the articulated kinematic structure of these actors remains underexplored. This leads us to our central question: Can we design a neural network for action learning that naturally integrates articulated kinematics as an inductive bias in its architecture?
To answer this question, we reflect on how CNNs [26] led to a deep learning framework that achieved remarkable success in computer vision by generalizing classical convolution filters from image processing [6]. Convolution filters naturally respect image structure through translation equivariance and spatial locality, but when used with handcrafted features, they were limited to low-level image processing tasks. Building on this foundation, CNNs extend these filters into neural operators with optimizable weights, non-linearities, and high-dimensional feature channels. This results in a deep learning framework capable of learning high-level semantic features from images while preserving the key structural properties of classical image filters.
@@
In a similar vein, we identify Rodrigues’ rotation formula as the basic operator in articulated forward kinematics and generalize it into a learnable form, which we call the Neural Rodrigues Operator. Concretely, we separate the entries in the Rodrigues’ rotation formula into state-dependent parameters conditioned on the joint angles, and state-independent coefficients that only rely on the articulated structure. We then extend it into a neural operator by treating the state-dependent parameters as input features, and relaxing the state-independent coefficients into optimizable weights. We further extend it into a multi-channel operator that applies to higher-dimensional features instead of 1D joint angles.
@@
With our neural operator as the key component, we construct a complete network architecture, named the Rodrigues Network (RodriNet), for encoding, understanding, and synthesizing articulated actions. It comprises three key components: a Rodrigues Layer for passing information from joints to links, built up our neural operator; a Joint Layer for passing information from links to joints; and finally, a Self-Attention Layer for global information exchange. We also introduce a global token for processing other task variables, such as perception inputs. Although starting with an operator in classical control theory, we end up with a deep neural network with modern designs while maintaining the structural bias in articulated kinematics, as illustrated in Figure 1.
@@
We evaluate our approach through three sets of experiments: First, we demonstrate the strong expressivity of our network in representing forward kinematics and motions through toy experiments. Second, we showcase our effectiveness in realistic robot-learning scenarios with imitation learning on 5 robot manipulation tasks, where our network achieves state-of-the-art results. Finally, we also show our network achieving state-of-the-art results in human hand pose estimation from images, where the articulated actor is no longer a robot, highlighting its applications in computer vision and graphics.
@@ 2 Related work: articulation-aware robot learning
Articulated robots can be naturally modeled as graphs, making graph convolution [4, 36] a popular approach for processing articulated data. This has been widely applied in character animation for tasks like skeleton-based action recognition [54, 13, 45, 46, 12, 5], pose estimation [16, 15, 55, 8], and motion retargeting [2]. Graph convolution effectively captures link connectivity and spatial locality, but it does not explicitly incorporate articulated kinematics, which is an essential aspect of articulated action understanding. In contrast, our work introduces a novel operator derived from forward kinematics, providing networks with kinematics-informed inductive bias.
Transformer-based architectures [48] are widely used in policy learning [3, 57, 9, 39, 43], and recent work has explored incorporating structural bias via graph-aware positional embeddings [22] or masked attention [41]. However, these modifications do not fundamentally adapt the self-attention mechanism to suit kinematics. We instead use standard self-attention layers for network capacity, but rely on our kinematics-inspired operator for inductive bias.
Forward kinematics provides a deterministic mapping from joint space to Cartesian space, but integrating it into neural networks remains non-trivial. Prior methods have inserted analytical forward kinematics as a differentiable layer in neural networks [49] to help them reason about the Cartesian results of the robot action. While this introduces kinematic awareness, it constrains the model’s flexibility. Other methods apply Cartesian-space loss functions after computing forward kinematics on network outputs [38, 23, 30], but these approaches do not alter the network architecture and are thus orthogonal to our focus. In contrast, our method derive a learnable operator from forward kinematics, thereby making the network kinematics-aware while maintaining the flexibility to learn high-level features.
@@ 3.1 Background
In this paper’s context, an articulated robot has a loop-free kinematic tree structure, with $D+1$ rigid links $L_{0},\cdots,L_{D}$ connected by $D$ one-DoF revolute joints $J_{1},\cdots,J_{D}$. This definition encompasses a wide range of platforms such as robotic arms, dexterous hands, quadrupeds, and humanoids. The kinematic tree has a root, denoted as the base link $L_{0}$, which can be either fixed or free-floating. All links and joints have their local 3D frames. Each joint $J_{j}$ connects a parent link $L_{\text{p}_{j}}$ with a child link $L_{\text{c}_{j}}$, where a rotation axis $\hat{\boldsymbol{\omega}}_{j}\in\mathbb{R}^{3}$ in the joint frame defines the rotational motion of the child link relative to the parent link, as well as a fixed transformation from the parent link’s frame to the joint’s frame $\mathbf{T}_{j}\in\mathrm{SE}(3)$.
Given the kinematic structure, the configuration of the robot is then determined by the joint angles $\boldsymbol{\theta}=[\theta_{1},\cdots,\theta_{D}]\in\mathbb{R}^{D}$, as well as the root pose $\mathbf{P}_{0}\in\mathrm{SE}(3)$ if its base link is free-floating (not required for fixed-base robots). In most classical control systems, control commands are sent to the joints, specifying their joint angles, velocities, or torques, depending on the control modes.
@@
where $\mathbf{t}_{i}\in\mathbb{R}^{3\times 1}$ is the position and $\mathbf{R}_{i}\in\mathbb{R}^{3\times 3}$ is the orientation. Given the pose of the base link $\mathbf{P}_{0}$, the poses of the descendant links can be computed recursively from parents to children. More concretely, at joint $J_{j}$, using $\mathbf{T}_{j},\hat{\boldsymbol{\omega}}_{j},\theta_{j}$, we can compute the child-link pose $\mathbf{P}_{\text{c}_{j}}$ from the parent-link pose $\mathbf{P}_{\text{p}_{j}}$ through two transformations: (i) a fixed coordinate change $\mathbf{T}_{j}$ from the parent frame to the joint frame; (ii) a dynamic rotation $\mathbf{R}(\hat{\boldsymbol{\omega}}_{j},\theta_{j})\in\mathbb{R}^{3\times 3}$ around axis $\hat{\boldsymbol{\omega}}_{j}$ of angle $\theta_{j}$ in the joint frame. Putting these together, the parent-to-children pose transformation is:
$\begin{bmatrix}\mathbf{R}_{\text{c}_{j}}&\mathbf{t}_{\text{c}_{j}}\\ \mathbf{0}_{1\times 3}&1\end{bmatrix}=\begin{bmatrix}\mathbf{R}_{\text{p}_{j}}&\mathbf{t}_{\text{p}_{j}}\\ \mathbf{0}_{1\times 3}&1\end{bmatrix}\mathbf{T}_{j}\begin{bmatrix}\mathbf{R}(\hat{\boldsymbol{\omega}}_{j},\theta_{j})&\mathbf{0}_{3\times 1}\\ \mathbf{0}_{1\times 3}&1\end{bmatrix}.$ (2)
Here, transformation (i) only depends on the robot’s articulated structure and thus is fixed, and transformation (ii) is state-dependent with variable $\theta_{j}$. Therefore, we can abbreviate Equation 2 as $\mathbf{P}_{\text{c}_{j}}=\mathbf{P}_{\text{p}_{j}}(\mathbf{T}_{j}\tilde{\mathbf{R}}(\hat{\boldsymbol{\omega}}_{j},\theta_{j}))$, where $\tilde{\mathbf{R}}(\hat{\boldsymbol{\omega}}_{j},\theta_{j})\in\mathbb{R}^{4\times 4}$ is the homogeneous matrix of the rotation. Essentially, forward kinematics is a hierarchical composition of fixed coordinate transformations and dynamic rotations in the axis-angle representation with variables $\theta_{1},\cdots,\theta_{D}$.
@@
The Rodrigues’ rotation formula tells how to compute the rotation matrix $\mathbf{R}(\hat{\boldsymbol{\omega}},\theta)\in\mathbb{R}^{3\times 3}$ from the axis-angle representation:
$\mathbf{R}(\hat{\boldsymbol{\omega}},\theta)=\mathbf{I}_{3}+\sin\theta[\hat{\boldsymbol{\omega}}]+(1-\cos\theta)[\hat{\boldsymbol{\omega}}]$ (3)
where $[\hat{\boldsymbol{\omega}}]\in\mathbb{R}^{3\times 3}$ is the skew-symmetric cross-product matrix of the rotation axis $\hat{\boldsymbol{\omega}}=(\hat{\omega}_{x},\hat{\omega}_{y},\hat{\omega}_{z})$.
@@ 3.2 Neural Rodrigues Operator with learnable parameters
Observe from Equation 3 that every entry in $\mathbf{R}(\hat{\boldsymbol{\omega}},\theta)$ is essentially a linear combination of $1$, $\cos\theta$, and $\sin\theta$, with fixed coefficients determined by the rotation axis $\hat{\boldsymbol{\omega}}$. Therefore, every entry in $\mathbf{P}(\hat{\boldsymbol{\omega}}_{j},\theta_{j})$ and thereby $\mathbf{T}_{j}\mathbf{P}(\hat{\boldsymbol{\omega}}_{j},\theta_{j})$ are linear combinations of $1$, $\cos\theta_{j}$, and $\sin\theta_{j}$, with constant coefficients defined by the state-independent parameters $\mathbf{T}_{j},\hat{\boldsymbol{\omega}}_{j}$ of joint $J_{j}$. Thus, we can re-parameterize Equation 2 as:
@@
where $\mathbf{A}_{j},\mathbf{B}_{j},\mathbf{C}_{j}\in\mathbb{R}^{4\times 4}$ are coefficient matrices that only depend on the robot’s articulated structure. Based on this, we construct our Neural Rodrigues Operator by replacing these fixed coefficients with learnable weights $W^{\text{bias}},W^{\text{cos}},W^{\text{sin}}\in\mathbb{R}^{4\times 4}$, resulting in:
@@ 3.3 Multi-Channel Neural Rodrigues Operator
Directly derived from the Rodrigues’ rotation formula, the neural operator defined in Equation 5 only applies to 1D joint features and $4\times 4$ link features. We now extend it to a multi-channel operator to learn higher-dimensional features. Specifically, we extend the link features from $4\times 4$ matrices to $F^{\text{in}}\in\mathbb{R}^{C_{L}\times 4\times 4},F^{\text{out}}\in\mathbb{R}^{C_{L}^{\prime}\times 4\times 4}$, and similarily, the joint features $\Theta\in\mathbb{R}^{C_{J}}$. Here $C_{L},C_{L}^{\prime},C_{J}$ are the channel numbers. Accordingly, the learnable weights are extended to $W^{\text{bias}}\in\mathbb{R}^{C_{L}\times C_{L}^{\prime}\times 4\times 4}$ and $W^{\text{cos}},W^{\text{sin}}\in\mathbb{R}^{C_{L}\times C_{L}^{\prime}\times C_{J}\times 4\times 4}$, giving a multi-channel extension of Equation 5:
@@ 4.1 Rodrigues Layer
With the Multi-Channel Rodrigues Operator being the core component operating on each joint, we construct the Rodrigues Layer, which extends this operator to the full tree structure of an articulated robot. We define the feature graph of an articulated robot as the collection of all link and joint features: $\{F_{l}\}_{j=0}^{D}$ for links and $\{\Theta_{j}\}_{j=1}^{D}$ for joints. The Rodrigues Layer maintains a set of Rodrigues Kernels $\{W_{j}^{*}\}_{j=1}^{D}$, one for each joint, and computes the output link features $\{F_{l}^{\text{out}}\}$ from the input link features $\{F_{l}^{\text{in}}\}$ and joint features $\{\Theta_{j}^{\text{in}}\}$. For each joint $J_{j}$, we retrieve its Rodrigues Kernels $W_{j}^{*}$, its joint feature $\Theta_{j}^{\text{in}}$, and the feature of its parent link $F^{\text{in}}_{\text{p}_{j}}$. We then apply the Multi-Channel Neural Rodrigues Operator (Equation 8) to compute the transformed feature:
@@
This transformed feature is passed to the joint’s child link $L_{\text{c}_{j}}$, added to the child’s input feature $F^{\text{in}}_{\text{c}_{j}}$, and normalized:
@@
For the root link without a parent, its output feature is defined as the layer-normalized version of its input. This formulation enables hierarchical, kinematics-aware feature propagation across the robot’s kinematic tree.

### v2

@@ Abstract
[…]
@@ 1 Introduction
We study the problem of understanding and predicting the actions of articulated actors. Articulated actors (e.g. robots (Shaw et al., 2023) and animated characters (Romero et al., 2022; Loper et al., 2023)) are systems that use multiple rotational joints to connect moving links. Their actions, including poses, motions, or control commands, are usually represented as values associated with all joints. Learning with articulated actors usually involves predicting their actions while processing diverse sensory inputs. Such a problem lies in a wide spectrum of intelligent systems, from whole-body controllers (Moro and Sentis, 2019; Kuang et al., 2025; Geng et al., 2025), to dexterous grasp detectors (Duan et al., 2021; Xu et al., 2023; Wan et al., 2023; Zhang et al., 2024), to motion retargeting networks (Aberman et al., 2020).
@@
These action data are inherently structured by articulated kinematics. This motivates us to design a neural network that leverages this structure as an inductive bias for better understanding and inference of actions. However, general architectures, such as MLPs and Transformers (Vaswani et al., 2017), treat actions as unstructured tokens, ignoring the kinematic relationships between joints. Previous works have exploited the connectivity of robot links through graph convolutions (Yan et al., 2018; Ci et al., 2019; Aberman et al., 2020) or masked attention (Sferrazza et al., 2024). Yet, architectures that directly exploit the computational patterns of articulated kinematics remain underexplored. This gap raises our central question: Can we design a neural network for action learning that embeds articulated kinematics as an inductive bias?
To answer this question, we draw an analogy to convolutional neural networks for images. Low-level 2D image features are spatially local and translation equivariant. Traditional pattern recognition exploits this structure using hand-crafted filters to detect edges (Canny, 1986) and corners (Harris et al., 1988). CNNs (LeCun et al., 2002) build on this idea by making filters learnable, adding non-linearities, and using high-dimensional channels. This creates a deep learning framework that learns high-level semantic features while preserving the structural properties of classical image filters.
@@
In a similar vein, we identify Rodrigues’ rotation formula as the fundamental operator in articulated forward kinematics and transform it into a learnable form, which we call the Neural Rodrigues Operator. Specifically, we separate the entries in the Rodrigues’ rotation formula into state-dependent parameters conditioned on the joint angles, and state-independent coefficients that only rely on the articulated structure. We then convert it into a neural operator by treating the state-dependent parameters as input features, and relaxing the state-independent coefficients into optimizable weights. We further extend it into a multi-channel operator that applies to higher-dimensional features rather than just 1D joint angles.
@@
With our neural operator as the key component, we construct a complete network architecture, named the Rodrigues Network (RodriNet), for encoding, understanding, and predicting articulated actions. It comprises three key components: a Rodrigues Layer for passing information from joints to links, built up our neural operator; a Joint Layer for passing information from links to joints; and finally, a Self-Attention Layer for global information exchange. We also introduce a global token for processing other task variables, such as perception inputs. Although starting with an operator in classical robotic theory, we end up with a deep neural network with modern designs while maintaining the structural bias in articulated kinematics, as illustrated in Figure 1.
@@
We evaluate our approach through three sets of experiments: First, we demonstrate the strong expressivity of our network in representing forward kinematics and motions through toy experiments. Second, we showcase our effectiveness in realistic robot-learning scenarios with imitation learning on 5 robot manipulation tasks. Additionally, we show our network achieving state-of-the-art results in human hand pose estimation from images, where the articulated actor is no longer a robot, highlighting its applications in computer vision and graphics.
@@ 2 Related work: articulation-aware robot learning
Articulated robots can be naturally modeled as graphs, making graph convolution (Bruna et al., 2013; Niepert et al., 2016) a popular approach for processing articulated data. This has been widely applied in character animation for tasks like skeleton-based action recognition (Yan et al., 2018; Cheng et al., 2020; Song et al., 2022; Chen et al., 2021b), pose estimation (Ci et al., 2019; Choi et al., 2020; Zeng et al., 2021), and motion retargeting (Aberman et al., 2020). Graph convolution effectively captures link connectivity and spatial locality, but it does not explicitly incorporate articulated kinematics, which is an essential aspect of articulated action understanding. In contrast, our work introduces a novel operator derived from forward kinematics, providing networks with kinematics-informed inductive bias.
Transformer-based architectures (Vaswani et al., 2017) are widely used in policy learning (Brohan et al., 2022; Zhao et al., 2023; Shridhar et al., 2023), and recent work has explored incorporating structural bias via graph-aware positional embeddings (Hong et al., 2021) or masked attention (Sferrazza et al., 2024). However, these modifications do not fundamentally adapt the self-attention mechanism to suit kinematics. We instead use standard self-attention layers for network capacity, but rely on our kinematics-inspired operator for inductive bias.
Forward kinematics provides a deterministic mapping from joint space to Cartesian space, but integrating it into neural networks remains non-trivial. Prior methods have inserted analytical forward kinematics as a differentiable layer in neural networks (Villegas et al., 2018) to help them reason about the Cartesian results of the robot action. While this introduces kinematic awareness, it constrains the model’s flexibility. Other methods apply Cartesian-space loss functions after computing forward kinematics on network outputs (Pavllo et al., 2020; Jiang et al., 2021; Liu et al., 2020), but these approaches do not alter the network architecture and are thus orthogonal to our focus. In contrast, our method derive a learnable operator from forward kinematics, thereby making the network kinematics-aware while maintaining the flexibility to learn high-level features.
@@ 3.1 Background
In this paper’s context, an articulated robot has a loop-free kinematic tree structure, with $D+1$ rigid links $L_{0},\cdots,L_{D}$ connected by $D$ one-DoF revolute joints $J_{1},\cdots,J_{D}$. This definition encompasses a wide range of platforms such as robotic arms, dexterous hands, quadrupeds, and humanoids. The kinematic tree has a root, denoted as the base link $L_{0}$, which can be either fixed or free-floating. All links and joints have their local 3D frames. Each joint $J_{j}$ connects a parent link $L_{\text{p}_{j}}$ with a child link $L_{\text{c}_{j}}$, where a rotation axis $\hat{\bm{\omega}}_{j}\in\mathbb{R}^{3}$ in the joint frame defines the rotational motion of the child link relative to the parent link, as well as a fixed transformation from the parent link’s frame to the joint’s frame $\mathbf{T}_{j}\in\mathrm{SE}(3)$.
Given the kinematic structure, the configuration of the robot is then determined by the joint angles $\bm{\theta}=[\theta_{1},\cdots,\theta_{D}]\in\mathbb{R}^{D}$, as well as the root pose $\mathbf{P}_{0}\in\mathrm{SE}(3)$ if its base link is free-floating (not required for fixed-base robots). In most classical control systems, control commands are sent to the joints, specifying their joint angles, velocities, or torques, depending on the control modes.
@@
where $\mathbf{t}_{i}\in\mathbb{R}^{3\times 1}$ is the position and $\mathbf{R}_{i}\in\mathbb{R}^{3\times 3}$ is the orientation. Given the pose of the base link $\mathbf{P}_{0}$, the poses of the descendant links can be computed recursively from parents to children. More concretely, at joint $J_{j}$, using $\mathbf{T}_{j},\hat{\bm{\omega}}_{j},\theta_{j}$, we can compute the child-link pose $\mathbf{P}_{\text{c}_{j}}$ from the parent-link pose $\mathbf{P}_{\text{p}_{j}}$ through two transformations: (i) a fixed coordinate change $\mathbf{T}_{j}$ from the parent frame to the joint frame; (ii) a dynamic rotation $\mathbf{R}(\hat{\bm{\omega}}_{j},\theta_{j})\in\mathbb{R}^{3\times 3}$ around axis $\hat{\bm{\omega}}_{j}$ of angle $\theta_{j}$ in the joint frame. Putting these together, the parent-to-children pose transformation is:
$\begin{bmatrix}\mathbf{R}_{\text{c}_{j}}&\mathbf{t}_{\text{c}_{j}}\\ \mathbf{0}_{1\times 3}&1\end{bmatrix}=\begin{bmatrix}\mathbf{R}_{\text{p}_{j}}&\mathbf{t}_{\text{p}_{j}}\\ \mathbf{0}_{1\times 3}&1\end{bmatrix}\mathbf{T}_{j}\begin{bmatrix}\mathbf{R}(\hat{\bm{\omega}}_{j},\theta_{j})&\mathbf{0}_{3\times 1}\\ \mathbf{0}_{1\times 3}&1\end{bmatrix}$. (2)
Here, transformation (i) only depends on the robot’s articulated structure and thus is fixed, and transformation (ii) is state-dependent with variable $\theta_{j}$. Therefore, we can abbreviate Equation 2 as $\mathbf{P}_{\text{c}_{j}}=\mathbf{P}_{\text{p}_{j}}(\mathbf{T}_{j}\tilde{\mathbf{R}}(\hat{\bm{\omega}}_{j},\theta_{j}))$, where $\tilde{\mathbf{R}}(\hat{\bm{\omega}}_{j},\theta_{j})\in\mathbb{R}^{4\times 4}$ is the homogeneous matrix of the rotation. Essentially, forward kinematics is a hierarchical composition of fixed coordinate transformations and dynamic rotations in the axis-angle representation with variables $\theta_{1},\cdots,\theta_{D}$.
@@
The Rodrigues’ rotation formula tells how to compute the rotation matrix $\mathbf{R}(\hat{\bm{\omega}},\theta)\in\mathbb{R}^{3\times 3}$ from the axis-angle representation:
$\mathbf{R}(\hat{\bm{\omega}},\theta)=\mathbf{I}_{3}+\sin\theta[\hat{\bm{\omega}}]+(1-\cos\theta)[\hat{\bm{\omega}}]^{2}$ (3)
where $\hat{\bm{\omega}}]\in\mathbb{R}^{3\times 3}$ is the skew-symmetric cross-product matrix of the rotation axis $\hat{\bm{\omega}}=(\hat{\omega}_{x},\hat{\omega}_{y},\hat{\omega}_{z})$.
@@ 3.2 Neural Rodrigues Operator with learnable parameters
Observe from Equation 3 that every entry in $\mathbf{R}(\hat{\bm{\omega}},\theta)$ is essentially a linear combination of $1$, $\cos\theta$, and $\sin\theta$, with fixed coefficients determined by the rotation axis $\hat{\bm{\omega}}$. Therefore, every entry in $\mathbf{P}(\hat{\bm{\omega}}_{j},\theta_{j})$ and thereby $\mathbf{T}_{j}\mathbf{P}(\hat{\bm{\omega}}_{j},\theta_{j})$ are linear combinations of $1$, $\cos\theta_{j}$, and $\sin\theta_{j}$, with constant coefficients defined by the state-independent parameters $\mathbf{T}_{j},\hat{\bm{\omega}}_{j}$ of joint $J_{j}$. Thus, we can re-parameterize Equation 2 as:
@@
where $\mathbf{A}_{j},\mathbf{B}_{j},\mathbf{C}_{j}\in\mathbb{R}^{4\times 4}$ are coefficient matrices that only depend on the robot’s articulated structure. Based on this, we construct our Neural Rodrigues Operator for one single joint by replacing these fixed coefficients with learnable weights $W^{\text{bias}},W^{\text{cos}},W^{\text{sin}}\in\mathbb{R}^{4\times 4}$, resulting in:
@@ 3.3 Multi-Channel Neural Rodrigues Operator
Derived from the Rodrigues’ rotation formula, the neural operator defined in Equation 5 only applies to 1D joint features and $4\times 4$ link features. We now extend it to a multi-channel operator to learn higher-dimensional features for a single joint. Specifically, we extend the link features from $4\times 4$ matrices to $F^{\text{in}}\in\mathbb{R}^{C_{L}\times 4\times 4},F^{\text{out}}\in\mathbb{R}^{C_{L}^{\prime}\times 4\times 4}$, and similarily, the joint features $\Theta\in\mathbb{R}^{C_{J}}$. Here $C_{L},C_{L}^{\prime},C_{J}$ are the channel numbers. Accordingly, the learnable weights are extended to $W^{\text{bias}}\in\mathbb{R}^{C_{L}\times C_{L}^{\prime}\times 4\times 4}$ and $W^{\text{cos}},W^{\text{sin}}\in\mathbb{R}^{C_{L}\times C_{L}^{\prime}\times C_{J}\times 4\times 4}$, giving a multi-channel extension of Equation 5:
@@ 4.1 Rodrigues Layer
With the Multi-Channel Rodrigues Operator being the core component operating on a single joint, we construct the Rodrigues Layer, which extends this operator to the full tree structure of an articulated robot. We define the feature graph of an articulated robot as the collection of all link and joint features: $\{F_{l}\}_{l=0}^{D}$ for links and $\{\Theta_{j}\}_{j=1}^{D}$ for joints. The Rodrigues Layer maintains a set of Rodrigues Kernels $\{W_{j}^{*}\}_{j=1}^{D}$, one for each joint, and computes the output link features $\{F_{l}^{\text{out}}\}$ from the input link features $\{F_{l}^{\text{in}}\}$ and joint features $\{\Theta_{j}^{\text{in}}\}$. For each joint $J_{j}$, we retrieve its Rodrigues Kernels $W_{j}^{*}$, its joint feature $\Theta_{j}^{\text{in}}$, and the feature of its parent link $F^{\text{in}}_{\text{p}_{j}}$. We then apply the Multi-Channel Neural Rodrigues Operator (Equation 8) to compute the transformed feature:
@@
This transformed feature is added to the child link $L_{\text{c}_{j}}$’s input feature $F^{\text{in}}_{\text{c}_{j}}$, and normalized:
@@
For the root link, its output feature is defined as its layer-normalized input. This layer updates the link features and leaves the joint features unchanged.